# Exact non-stationary solutions of the Euler equations in two and three dimensions

### v1

@@ 4. Three-Dimensional Fluids
We now consider three-dimensional manifolds, for which the inertia operator is $\mathcal{A}=\operatorname{curl}$. To construct examples as in Theorem 2.2, we must understand when a manifold admits a Killing field $X$ whose vorticity $\operatorname{curl}X$ is also Killing. As we will see, the analysis naturally splits into two cases depending on whether $\lvert X\rvert_{g}$ is constant. Independently of the Killing assumption, Arnold’s structure theorem for steady Euler flows [2, Section II.1] implies that away from singular sets either $\operatorname{curl}X=fX$ for some function $f$ or the flow lines lie on invariant two-dimensional tori. When both $X$ and $\operatorname{curl}X$ are Killing, this dichotomy becomes geometrically rigid. The following lemma provides the key rigidity input.
Lemma 4.1.
@@
Proof.
(4.1) $g(\nabla_{v}X,w)+g(v,\nabla_{w}X)=0$
and
(4.2) $g(\nabla_{v}\operatorname{curl}X,w)+g(v,\nabla_{w}\operatorname{curl}X)=0$
respectively, for all $v,w\in\mathfrak{X}(M)$. Hence, letting $w=X$ in (4.2) yields that, for all $v\in\mathfrak{X}(M)$, we have
$g(\nabla_{v}\operatorname{curl}X,X)+g(v,\nabla_{X}\operatorname{curl}X)=0.$
Furthermore, we have [^3] See [20] for example.
(4.3) $X\times\operatorname{curl}{X}=\nabla\lvert X\rvert_{g}^{2}$
In particular $X$ and $\operatorname{curl}X$ are linearly independent on an open set if and only if $\lvert X\rvert_{g}$ is non-constant. Taking the curl of the left hand side of (4.3) gives
$\operatorname{curl}\big(X\times\operatorname{curl}X\big)$ $=\nabla_{\operatorname{curl}X}X-\nabla_{X}\operatorname{curl}X+\mathrm{div}(\operatorname{curl}X)X-\mathrm{div}(X)\operatorname{curl}X$
$=\nabla_{\operatorname{curl}X}X-\nabla_{X}\operatorname{curl}X.$
Hence, as $\operatorname{curl}\nabla\lvert X\rvert_{g}^{2}=0$, we have $\nabla_{X}\operatorname{curl}X=\nabla_{\operatorname{curl}X}X$.
Now, using the above, for any $v\in\mathfrak{X}(M)$ we have
$v\big(g(X,\operatorname{curl}X)\big)$ $=g(\nabla_{v}X,\operatorname{curl}X)+g(X,\nabla_{v}\operatorname{curl}X)$
$=g(\nabla_{v}X,\operatorname{curl}X)-g(\nabla_{X}\operatorname{curl}X,v)$
$=g(\nabla_{v}X,\operatorname{curl}X)-g(\nabla_{\operatorname{curl}X}X,v).$
By definition of $\operatorname{curl}$ and the cross product we have
$g(\nabla_{v}X,\operatorname{curl}X)=g(\nabla_{\operatorname{curl}X}X,v)+g(\operatorname{curl}{X},v\times\operatorname{curl}X)=g(\nabla_{\operatorname{curl}X}X,v).$
Putting all of this together gives
$v\big(g(X,\operatorname{curl}X)\big)=0.$
Hence $g(X,\operatorname{curl}X)$ is constant on $M$.
Lastly, if $\operatorname{curl}X=fX$ for some function $f$, then by (4.3), we must have that $\lvert X\rvert_{g}$ is constant, and hence $f=\frac{g(X,\operatorname{curl}X)}{\lvert X\rvert_{g}^{2}}$ is also constant. ∎
Hence, Lemma 4.1 reduces our analysis to two geometrically distinct situations.
If $\operatorname{curl}X$ is everywhere parallel to $X$, then $\operatorname{curl}X=\alpha X$ for some constant $\alpha$, and $X$ has constant length. This is the strong Beltrami case, studied in Section 4.1, where the manifold is a circle bundle over a surface, whose geometry is largely unrestricted.
@@ 4.1. Circle Bundles over a Surface
Here we consider the situation where $X$ is a Killing field on a compact orientable three-dimensional manifold $(M,g)$ for which $\operatorname{curl}{X}$ is also a Killing field that is that is everywhere parallel to $X$. Lemma 4.1 shows that in fact $X$ must be nowhere zero and of constant length. It is possible that such a Killing field could be irregular, such as $X=a\,\partial_{x}+b\,\partial_{y}+c\,\partial_{z}$ on the flat torus $\mathbb{T}^{3}$, with $\mathcal{A}X=0$; if the coefficient ratios are irrational, the orbits of $X$ can be dense.
The structure theory is simpler if we assume $X$ is regular, with closed orbits all of the same length, so that we have a circle bundle over a compact orientable surface $\Sigma$. All such manifolds have a well-known classification due originally to Boothby-Wang [6], uniquely characterized up to bundle isomorphism by the genus of the surface (a nonnegative integer) and the Euler class (any integer), and for each choice of integers there is a unique compact orientable $M$ which is a circle bundle over it with that Euler class.
Furthermore the fact that $X$ is a nowhere-zero curl eigenfield implies that its dual $1$-form $X^{\flat}$ is a contact form, and since $X$ is also Killing, we get a K-contact structure, which in three dimensions implies that $M$ is a Sasakian manifold. Using the classification of Sasakian metrics [5], we see that all the topological and metric data on $M$ is determined by an arbitrary Riemannian metric on the closed orientable surface $\Sigma$ and the Euler class of the bundle.
If the Euler class is zero, the bundle is trivial, and $M=\Sigma\times S^{1}$ with the product Riemannian metric. In this case $X$ is a parallel field (since it is both Killing and harmonic), so that $\lambda=0$ in Theorem 2.2 and the time dependence is essentially trivial (the action by translation in the $S^{1}$ variable which does not twist anything). To get genuinely nontrivial time dependence, we need nonzero Euler class; fortunately there are many examples, including most famously the Hopf fibration of $S^{3}$ over $S^{2}$ (discussed in Example 4.5).
We will first describe the general classification of circle bundles over compact manifolds admitting Killing fields that are also curl eigenfields in Section 4.1, showing that for an arbitrary Riemann surface $\Sigma$ and an arbitrary nonzero integer, there is a circle bundle $M$ with this structure. In Section 4.1.2, we extend the Chandrasekhar-Kendall construction to show how to obtain the curl eigenfields on any $3$-manifold with this structure, in terms of eigenfunctions of the Laplacian on $M$. Finally in Example 4.5 we work out all the details for the Hopf fibration of $S^{3}$ over $S^{2}$, including explicit formulas for the exact non-steady solutions.
@@ 4.1.1. Geometric theory of circle bundles
Conversely given a compact oriented surface $\Sigma$ with any Riemannian metric and any nonzero integer $k$, there is a unique (up to isomorphism) compact oriented $3$-manifold $M$, a map $\tau\colon M\to\Sigma$, and a contact form $\omega$ on $M$ such that $\tau$ is a Riemannian submersion, $\omega$ is the connection $1$-form of the bundle, the Reeb field $X=\omega^{\sharp}$ is unit-length and Killing with orbits of length $2\pi$, and $\operatorname{curl}{X}=2\epsilon X$ for $\epsilon=\pi k/\mathrm{Area}(\Sigma)$.
@@
Example 4.3.
Let
(4.4) $s_{k}(r):=\begin{cases}\sin{r}&k=1,\\ r&k=0,\\ \sinh{r}&k=-1;\end{cases}\qquad\text{and}\qquad c_{k}(r):=\begin{cases}\cos{r}&k=1,\\ 1&k=0,\\ \cosh{r}&k=-1.\end{cases}$
Choose $\varphi$ so that $\varphi_{r}(r,\psi)=2s_{k}(r/2)^{2}$, which makes $\varphi_{rr}(r,\psi)=s_{k}(r)$. Then a metric satisfying these conditions is given by
$ds^{2}=\big(d\theta+2s_{k}(r/2)^{2}\,d\psi\big)^{2}+dr^{2}+s_{k}(r)^{2}\,d\psi^{2}.$
@@
More generally we could choose geodesic polar coordinates on $\Sigma$ so that the metric is $g_{\Sigma}=dr^{2}+h(r,\psi)^{2}\,d\psi^{2}.$ Then the area form is $dA_{\Sigma}=h(r,\psi)\,dr\wedge d\psi$, and a $1$-form $\eta$ on $M$ with the desired properties would be
$\eta=d\theta-2\epsilon f_{\psi}\,dr+2\epsilon f_{r}\,d\psi$
with $f$ solving the Laplace equation $f_{rr}+f_{\psi\psi}=h(r,\psi)$. The metric would then take the Kaluza-Klein form
$ds^{2}=\big(d\theta-2\epsilon f_{\psi}\,dr+2\epsilon f_{r}\,d\psi\big)^{2}+dr^{2}+(f_{rr}+f_{\psi\psi})^{2}\,d\psi^{2},$
@@ 4.1.2. Curl eigenfields in a circle bundle
(4.5) $\operatorname{curl}X,z]=i\lambda\operatorname{curl}z,\qquad\operatorname{curl}z=\alpha z,\qquad[X,z]=i\zeta z,$
@@
(4.6) $\operatorname{curl}z=\alpha z,\qquad[X,z]=inz,$
@@
The following is a generalization of the Chandrasekhar-Kendall [9] construction of curl eigenfields, which reduces the problem to finding eigenfunctions of the Laplacian. It may be viewed as a converse to Proposition 1 in [27], where it is shown that (in our notation) if $z$ is a curl eigenfield with eigenvalue $\alpha$, then $g(X,z)$ is a Laplacian eigenfunction with eigenvalue $\delta^{2}=\alpha(\alpha-2\epsilon)$.
@@
Suppose $(M,g)$ is a three-dimensional Riemannian manifold, and that $X$ is a Killing field on $M$. Assume $X$ is a curl eigenfield with $\operatorname{curl}{X}=2\epsilon X$ for some $\epsilon\in\mathbb{R}$. Let $f$ be a complex-valued function on $M$ which satisfies
(4.7) $\Delta f=-\delta^{2}f,\qquad X(f)=inf.$
for a real number $\delta>0$ and an integer $n$. Then for $\alpha_{\pm}:=\epsilon\pm\sqrt{\epsilon^{2}+\delta^{2}}$, the vector fields
$z_{\pm}=\alpha_{\pm}^{2}fX+\alpha_{\pm}\nabla f\times X+in\nabla f$
are curl eigenfields on $M$ with eigenvalues $\alpha_{\pm}$. Any such $z_{\pm}$ gives an exact solution of the $3D$ Euler equation as in Theorem 2.2.
@@
Existence of $f$ satisfying (4.7) comes from the fact that $X$ is a Killing field, so it commutes with the Laplacian, and we can find simultaneous eigenvectors of these operators.
Define vector fields
(4.8) $S_{1}:=fX,\qquad S_{2}:=\nabla f\times X,\qquad S_{3}:=\nabla f.$
Then curl maps these fields to constant linear combinations of themselves, via
$\operatorname{curl}(S_{1})$ $=\nabla f\times X+f\operatorname{curl}{X}=2\epsilon S_{1}+S_{2}$
$\operatorname{curl}(S_{2})$ $=(\operatorname{div}{X})\nabla f-(\Delta f)X+[X,\nabla f]=\delta^{2}fX+in\nabla f=\delta^{2}S_{1}+inS_{3}$
$\operatorname{curl}(S_{3})$ $=0.$
@@
Here we used the fact that $\operatorname{div}{X}=0$ and that, since $X$ is a Killing field, its action on functions commutes with the gradient, so that $X,\nabla f]=\nabla X(f)$.
Thus curl on the span of these vector fields is given by the $3\times 3$ matrix

### v2

@@ 4. Three-Dimensional Fluids
We now consider three-dimensional manifolds, for which the inertia operator is $\mathcal{A}=\operatorname{curl}$, cf. (2.3). To construct examples as in Theorem 2.2, we must understand when a manifold admits a Killing field $X$ whose vorticity $\operatorname{curl}X$ is also Killing. As we will see, the analysis naturally splits into two cases depending on whether $\lvert X\rvert_{g}$ is constant. Independently of the Killing assumption, Arnold’s structure theorem for steady Euler flows [2, Section II.1] implies that away from singular sets either $\operatorname{curl}X=fX$ for some function $f$ or the flow lines lie on invariant two-dimensional tori. When both $X$ and $\operatorname{curl}X$ are Killing, this dichotomy becomes geometrically rigid. The following lemma provides the key rigidity input.
Proposition 4.1.
@@
While the proof of the above can be carried out locally in coordinates, a global formulation requires intrinsic, coordinate-free computations involving the Levi-Civita connection. We defer this to the Appendix.
Hence, as mentioned above, Proposition 4.1 reduces our analysis to two geometrically distinct situations.
If $\operatorname{curl}X$ is everywhere parallel to $X$, then $\operatorname{curl}X=\alpha X$ for some constant $\alpha$, and $X$ has constant length. This is the strong Beltrami case, studied in Section 4.1, where the manifold is a circle bundle over a Riemann surface with arbitrary geometry.
@@ 4.1. Circle Bundles over a Surface
Here we consider the situation where $X$ is a Killing field on a compact orientable three-dimensional manifold $(M,g)$ for which $\operatorname{curl}{X}$ is also a Killing field that is everywhere parallel to $X$. Proposition 4.1 shows that in fact $X$ must be nowhere zero and of constant length. It is possible that such a Killing field could be irregular, such as $X=a\,\partial_{x}+b\,\partial_{y}+c\,\partial_{z}$ on the flat torus $\mathbb{T}^{3}$, with $\mathcal{A}X=0$; if the coefficient ratios are irrational, the orbits of $X$ can be dense. The structure theory is simpler if we assume $X$ is regular, with closed orbits all of the same length, so that we have a circle bundle over a compact orientable surface $\Sigma$ with some integer Euler class.
If the Euler class is zero, the bundle is trivial, and $M=\Sigma\times S^{1}$ with the product Riemannian metric. In this case $X$ is a parallel field (since it is both Killing and harmonic), so that $\lambda=0$ in Theorem 2.2 and the time dependence is essentially trivial (the action by translation in the $S^{1}$ variable which does not twist anything, as in Remark 2.4). To get genuinely nontrivial time dependence, we need nonzero Euler class.
We will first describe the general classification of circle bundles over compact manifolds admitting Killing fields that are also curl eigenfields in Section 4.1. Using tools from contact geometry we show that for an arbitrary Riemann surface $\Sigma$ and an arbitrary nonzero integer, there is a circle bundle $M$ with this structure. In Section 4.1.2, we begin with Example 4.4 where we consider the Hopf-fibration of $S^{3}$ over $S^{2}$ and give explicit formulas for exact non-steady solutions of (1.1) in this setting. We then extend the Chandrasekhar-Kendall construction of curl eigenfields on any three-dimensional manifold with this structure, in terms of eigenfunctions of the Laplacian on $M$. We conclude by discussing the flat and negatively curved analogues.
@@ 4.1.1. Geometric theory of circle bundles
Conversely given a compact oriented surface $\Sigma$ with any Riemannian metric and any nonzero integer $k$, there is a compact oriented $3$-manifold $M$, a map $\tau\colon M\to\Sigma$, and a contact form $\eta$ on $M$ such that $\tau$ is a Riemannian submersion, $\omega$ is the connection $1$-form of the bundle, the Reeb field $X=\eta^{\sharp}$ is unit-length and Killing with orbits of length $2\pi$, and $\operatorname{curl}{X}=2\epsilon X$ for $\epsilon=\pi k/\mathrm{Area}(\Sigma)$.
@@
Remark 4.3.
More explicitly, we can choose geodesic polar coordinates on $\Sigma$ so that the metric is given by $g_{\Sigma}=dr^{2}+h(r,\sigma)^{2}\,d\sigma^{2}.$ Then the area form is $dA_{\Sigma}=h(r,\sigma)\,dr\wedge d\sigma$, and a $1$-form $\eta$ on $M$ with the desired properties is given locally by
$\eta=d\theta-2\epsilon f_{\sigma}\,dr+2\epsilon f_{r}\,d\sigma$
with $f$ solving the Laplace equation $f_{rr}+f_{\sigma\sigma}=h(r,\sigma)$. The metric then takes the Kaluza-Klein form
$ds^{2}=\big(d\theta-2\epsilon f_{\sigma}\,dr+2\epsilon f_{r}\,d\sigma\big)^{2}+dr^{2}+(f_{rr}+f_{\sigma\sigma})^{2}\,d\sigma^{2},$
For example, choosing $\varphi$ so that $\varphi_{r}(r,\sigma)=2s_{\kappa}(\tfrac{r}{2})^{2}$, makes $\varphi_{rr}(r,\sigma)=s_{\kappa}(r)$, in terms of the generalized trigonometric functions $s_{\kappa}$ from (3.1). The metric is then given locally by
$ds^{2}=\big(d\theta+2s_{\kappa}(\tfrac{r}{2})^{2}\,d\sigma\big)^{2}+dr^{2}+s_{\kappa}(r)^{2}\,d\sigma^{2}.$
@@ 4.1.2. Curl eigenfields in a circle bundle
(4.1) $\operatorname{curl}X,\mathfrak{q}]=i\lambda\operatorname{curl}\mathfrak{q},\qquad\operatorname{curl}\mathfrak{q}=\alpha\mathfrak{q},\qquad[X,\mathfrak{q}]=i\zeta\mathfrak{q},$
@@
(4.2) $\operatorname{curl}\mathfrak{q}=\alpha\mathfrak{q},\qquad[X,\mathfrak{q}]=in\mathfrak{q},$
@@
We begin with the case of the round three-sphere, where a convenient basis of curl eigenfields is already known.
The Hopf fibration is the most well-known nontrivial circle bundle, mapping $S^{3}$ to $S^{2}$. Given the Hopf fields
$e_{1}$ $=-y\partial_{x}+x\partial_{y}-w\partial_{z}+z\partial_{w},$
$e_{2}$ $=-z\partial_{x}+w\partial_{y}+x\partial_{z}-y\partial_{w},$
$e_{3}$ $=-w\partial_{x}-z\partial_{y}+y\partial_{z}+x\partial_{w}.$
we borrow a basis of curl eigenfields from [5, Appendix A] which is convenient for our purposes. In particular,
$\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(S^{3})=\bigcup_{k=0}^{\infty}\bigcup_{m=0}^{k}\mathcal{E}_{k}^{m}\ \cup\ \bigcup_{k=2}^{\infty}\bigcup_{m=1}^{k-1}\mathcal{F}_{k}^{m}$
such that for $k\geq 1$ if $\mathfrak{q}\in\mathcal{E}_{k}^{m}$ then $\operatorname{curl}\mathfrak{q}=(k+2)v$ and $\mathfrak{q},e_{1}]=-i(2m-k)\mathfrak{q}$. Hence, for $\mathfrak{q}\in\mathcal{E}_{k}^{m}$, in the context of Theorem 2.2, letting $u_{0}=e_{1}$, we have
$K_{e_{1}}\mathfrak{q}=-i\frac{2(2m-k)}{k+2}\mathfrak{q},\qquad\operatorname{curl}\mathfrak{q}=(k+2)\mathfrak{q},\qquad[e_{1},\mathfrak{q}]=i(2m-k)\mathfrak{q}.$
Note that corresponding solutions of the Euler equations generated by these basis elements are stationary [^5] The solutions generated by the elements of $\mathcal{E}_{0}^{0}$ are all stationary. if and only if $\varpi_{k,m}=(2m-k)\left(\frac{2}{k+2}-1\right)=0$. That is, if $m=\frac{k}{2}$. Hence, the simplest non-stationary solutions come from setting $k=1$ and $m=0$, which gives $\varpi_{1,0}=\frac{1}{3}$. We briefly recall the construction of the basis.
Identifying $\mathbb{R}^{4}\simeq\mathbb{C}^{2}$ we introduce complex coordinates
$\alpha=x+iy,\qquad\bar{\alpha}=x-iy,\qquad\beta=z+iw,\qquad\bar{\beta}=z-iw$
and define the family of homogeneous polynomials in the formal variables $z_{1},z_{2}$
$F_{k}^{m}(\alpha,\bar{\alpha},\beta,\bar{\beta})(z_{1},z_{2})=(\alpha z_{1}+\beta z_{2})^{m}(-\bar{\beta}z_{1}+\bar{\alpha}z_{2})^{k-m},$
for integer $k\geq 0$ and $0\leq m\leq k$ along with, for $0\leq j\leq k$, the implicitly defined $Q^{m}_{kj}(\alpha,\bar{\alpha},\beta,\bar{\beta})$ given by
(4.3) $F_{k}^{m}(\alpha,\bar{\alpha},\beta,\bar{\beta})(z_{1},z_{2})=\sum_{j=0}^{k}Q^{m}_{kj}(\alpha,\bar{\alpha},\beta,\bar{\beta})z_{1}^{j}z_{2}^{k-j}.$
The $Q^{m}_{kj}$ are eigenfunctions of the Laplacian on $S^{3}$ with eigenvalues $k(k+2)$.
The simplest basis element [^6] See [5, Appendix A] for the definitions of the basis elements. from $\mathcal{E}_{1}^{0}$ is
$v_{112}^{0}=Q_{11}^{0}(e_{2}-ie_{3})=(x-iy)(e_{2}-ie_{3})$
and the corresponding real solution is the restriction to $S^{3}$ of
(4.4) $U_{\mathfrak{R}}(t)=e_{1}+\left(xe_{2}-ye_{3}\right)\cos\left(\frac{t}{3}\right)+\left(ye_{2}+xe_{3}\right)\sin\left(\frac{t}{3}\right)$
which, as noted before, is of a similar spirit to the two-dimensional Rossby-Haurwitz waves given in (3.4).
We now outline a recipe for constructing fields on manifolds of the type described in Theorem 4.2 satisfying the conditions (4.2)—and thus the assumptions of Theorem 2.2—from eigenfunctions of the Laplacian. It is a generalization of the Chandrasekhar-Kendall [10] construction and may be viewed as a converse to Proposition 1 in [30], where it is shown that (in our notation) if $\mathfrak{q}$ is a curl eigenfield with eigenvalue $\alpha$, then $g(X,\mathfrak{q})$ is a Laplacian eigenfunction with eigenvalue $\delta^{2}=\alpha(\alpha-2\epsilon)$.
@@
Let $(M,g)$ be a three-dimensional Riemannian manifold which admits a Killing field $X$ such that $\operatorname{curl}{X}=2\epsilon X$ for some $\epsilon\in\mathbb{R}$. If $f$ is a complex-valued function on $M$ with constants $\delta>0$ and $\zeta\in\mathbb{R}$ such that
(4.5) $\Delta f=\delta^{2}f,\qquad X(f)=i\zeta f,$
then, letting $\alpha_{\pm}:=\epsilon\pm\sqrt{\epsilon^{2}+\delta^{2}}$, the vector fields
$\mathfrak{q}_{\pm}=\alpha_{\pm}^{2}fX+\alpha_{\pm}\nabla f\times X+i\zeta\nabla f$
are curl eigenfields on $M$ with eigenvalues $\alpha_{\pm}$.
Furthermore, any such $\mathfrak{q}_{\pm}$ yields an exact solution of the Euler equations (1.1) via Theorem 2.2.
@@
For any $f$ satisfying the assumptions of (4.5), consider the vector fields
(4.6) $E_{1}=fX,\qquad E_{2}=\nabla f\times X,\qquad E_{3}=\nabla f.$
Taking the $\operatorname{curl}$ of these fields yields
$\operatorname{curl}(E_{1})$ $=\nabla f\times X+f\operatorname{curl}{X}=2\epsilon E_{1}+E_{2}$
$\operatorname{curl}(E_{2})$ $=(\operatorname{div}{X})\nabla f+(\Delta f)X+[X,\nabla f]=\delta^{2}fX+i\zeta\nabla f=\delta^{2}E_{1}+i\zeta E_{3}$
$\operatorname{curl}(E_{3})$ $=0.$
@@
where we have used the fact that $\operatorname{div}{X}=0$ and that, since $X$ is a Killing field, its action on functions commutes with the gradient, giving $X,\nabla f]=\nabla X(f)$. Hence curl on the span of these vector fields is given by the matrix